MONT 100N – Modeling the Environment Purgatory Chasm Trip – Saturday, October 21

Note: Some information about the trip has changed – updated details:

1. Departure from HC will be 10:00am sharp from Hogan 3 (not Kimball); Be there by 9:45am for safety.

2. Departure from Purgatory Chasm will be at 1:00pm sharp (we’ll be back to campus earlier than previously planned).

3. When you get on the bus, I will check off your name on a class roster and you will get a randomly-assigned ticket for the geology walks with Prof. Sara Mitchell:

a. Group 1: 10:30 - 10:45 b. Group 2: 10:55 - 11:10 c. Group 3: 11:20 - 11:35 d. Group 4: 11:45 - 12:00 e. Group 5: 12:10 - 12:25

We will give instructions about where to meet your group when we arrive at the park.

4. Other than the time when you are doing the geology walk, you are free to explore the trails, sketch leaves, determine information for your papers, etc. (See map on reverse for hiking trails in the park.)

5. Because of the closure of Kimball on Saturday for the alumni events, we were only able to get snacks/fruit/drinks, not a whole lunch. You may want to eat before we leave. If not, you should still be able to get lunch on campus when we return.

Suggestions and Guidelines

To make sure everyone has a safe and enjoyable experience, know that you will be hiking over a trail that includes slippery boulders worn smooth by other hikers in a (somewhat) natural setting,

• wear sturdy shoes – sneakers, hiking boots, etc. NO SANDALS, “FLIP-FLOPS” etc, PLEASE!

• Stay away from the edges of the chasm and do not climb the chasm walls or large boulders

• the weather forecast looks ideal as of now – low 70’s Fahrenheit and clear/sunny, but be aware that things can change

• wearing sunscreen recommended (Prof. Little will bring some)

• wearing bug spray against ticks, mosquitos, etc. also recommended (Prof. Little will bring some).

• bring a backpack to carry notebook, paper, writing implements for notes, etc.

• bring your phone if you want to take pictures (but also try to experience some of the scenery “first-hand” without being too tied to your screen!)